# On the torsion growth in quadratic number fields for elliptic curves defined over the rationals

### v1

@@ A Tate normal form is an equation like
So, $v_{2}(x_{1})=1$ and $v_{2}(b_{2})=0$.
Let us now bound the valuation of
$s=\frac{8b_{4}+2b_{2}x_{1}+3x_{1}^{2}}{2y_{1}}.$
First of all, note that the numerator is the sum of a term with $2$-adic valuation at least $3$ and two terms with $2$-adic valuation exactly $2$. Since the denominator has $2$-adic valuation equal to $2$, we obtain that $v_{2}(s)\geq 0$. In fact, we can even argue that $v_{2}(s)\geq 1$, since the sum of two terms of $2$-adic valuation equal to $2$ will have valuation greater than or equal to $3$.
Summing up, we have the following information on the valuations of our constants:
$v_{2}(q)=2,v_{2}(y_{1})=v_{2}(x_{1})=1,v_{2}(b_{2})=0,v_{2}(s)\geq 1.$
From Equation (14),
$2v_{2}(q)+8=4v_{2}(y_{1})+v_{2}(q+4p)+v_{2}(q-4p),$
and we can deduce
$4=v_{2}(1+16b)=v_{2}(\overline{a}_{3}^{2}+16\overline{a}_{2}^{3})-4,$
which implies that
$8=v_{2}\left(4y_{1}^{2}+16(-s^{2}+3x_{1}+b_{2})^{3}\right),$
where (friendly remainder),
$s=\frac{8b_{4}+2b_{2}x_{1}+3x_{1}^{2}}{2y_{1}},\qquad y_{1}^{2}=x_{1}^{3}+b_{2}x_{1}^{2}+8b_{4}x_{1}+16b_{6}.$
Now we substitute $y_{1}^{2}$ in $v_{2}$ and then $s$. That is we only rewrite $y_{1}^{2}$ as a polynomial in $x_{1}$ once, in the term $4y_{1}^{2}$. This way we obtain the (rather cumbersome) expression:
88 88 $=$ $-v_{2}(-4y_{1}^{6})+v_{2}\Big{(}64b_{2}^{6}x_{1}^{6}+576b_{2}^{5}x_{1}^{7}+2160b_{2}^{4}x_{1}^{8}+4320b_{2}^{3}x_{1}^{9}+4860b_{2}^{2}x_{1}^{10}+2916b_{2}x_{1}^{11}+$
$\quad 729x_{1}^{12}+1536b_{2}^{5}b_{4}x_{1}^{5}+11520b_{2}^{4}b_{4}x_{1}^{6}+34560b_{2}^{3}b_{4}x_{1}^{7}+51840b_{2}^{2}b_{4}x_{1}^{8}+38880b_{2}b_{4}x_{1}^{9}+$
$\quad 11664b_{4}x_{1}^{10}-192b_{2}^{5}x_{1}^{4}y_{1}^{2}-1728b_{2}^{4}x_{1}^{5}y_{1}^{2}-6048b_{2}^{3}x_{1}^{6}y_{1}^{2}-10368b_{2}^{2}x_{1}^{7}y_{1}^{2}-8748b_{2}x_{1}^{8}y_{1}^{2}-$
$\quad 2916x_{1}^{9}y_{1}^{2}+15360b_{2}^{4}b_{4}^{2}x_{1}^{4}+92160b_{2}^{3}b_{4}^{2}x_{1}^{5}+207360b_{2}^{2}b_{4}^{2}x_{1}^{6}+207360b_{2}b_{4}^{2}x_{1}^{7}+$
$\quad 77760b_{4}^{2}x_{1}^{8}-3072b_{2}^{4}b_{4}x_{1}^{3}y_{1}^{2}-23040b_{2}^{3}b_{4}x_{1}^{4}y_{1}^{2}-62208b_{2}^{2}b_{4}x_{1}^{5}y_{1}^{2}-72576b_{2}b_{4}x_{1}^{6}y_{1}^{2}-$
$\quad 31104b_{4}x_{1}^{7}y_{1}^{2}+192b_{2}^{4}x_{1}^{2}y_{1}^{4}+1728b_{2}^{3}x_{1}^{3}y_{1}^{4}+5616b_{2}^{2}x_{1}^{4}y_{1}^{4}+7776b_{2}x_{1}^{5}y_{1}^{4}+3888x_{1}^{6}y_{1}^{4}+$
$\quad 81920b_{2}^{3}b_{4}^{3}x_{1}^{3}+368640b_{2}^{2}b_{4}^{3}x_{1}^{4}+552960b_{2}b_{4}^{3}x_{1}^{5}+276480b_{4}^{3}x_{1}^{6}-18432b_{2}^{3}b_{4}^{2}x_{1}^{2}y_{1}^{2}-$
$\quad 110592b_{2}^{2}b_{4}^{2}x_{1}^{3}y_{1}^{2}-207360b_{2}b_{4}^{2}x_{1}^{4}y_{1}^{2}-124416b_{4}^{2}x_{1}^{5}y_{1}^{2}+1536b_{2}^{3}b_{4}x_{1}y_{1}^{4}+$
$\quad 11520b_{2}^{2}b_{4}x_{1}^{2}y_{1}^{4}+27648b_{2}b_{4}x_{1}^{3}y_{1}^{4}+20736b_{4}x_{1}^{4}y_{1}^{4}-64b_{2}^{3}y_{1}^{6}-576b_{2}^{2}x_{1}y_{1}^{6}-$
$\quad 1744b_{2}x_{1}^{2}y_{1}^{6}-1744x_{1}^{3}y_{1}^{6}+245760b_{2}^{2}b_{4}^{4}x_{1}^{2}+737280b_{2}b_{4}^{4}x_{1}^{3}+552960b_{4}^{4}x_{1}^{4}-$
$\quad 49152b_{2}^{2}b_{4}^{3}x_{1}y_{1}^{2}-221184b_{2}b_{4}^{3}x_{1}^{2}y_{1}^{2}-221184b_{4}^{3}x_{1}^{3}y_{1}^{2}+3072b_{2}^{2}b_{4}^{2}y_{1}^{4}+18432b_{2}b_{4}^{2}x_{1}y_{1}^{4}+$
$\quad 27648b_{4}^{2}x_{1}^{2}y_{1}^{4}-128b_{4}x_{1}y_{1}^{6}+393216b_{2}b_{4}^{5}x_{1}+589824b_{4}^{5}x_{1}^{2}-49152b_{2}b_{4}^{4}y_{1}^{2}-$
$\quad 147456b_{4}^{4}x_{1}y_{1}^{2}-256b_{6}y_{1}^{6}+262144b_{4}^{6}\Big{)}$
As we know $v_{2}(y_{1})=1$, we have
1616 1616 $=$ $v_{2}\Big{(}64b_{2}^{6}x_{1}^{6}+576b_{2}^{5}x_{1}^{7}+2160b_{2}^{4}x_{1}^{8}+4320b_{2}^{3}x_{1}^{9}+4860b_{2}^{2}x_{1}^{10}+2916b_{2}x_{1}^{11}+729x_{1}^{12}+$
$\quad 1536b_{2}^{5}b_{4}x_{1}^{5}+11520b_{2}^{4}b_{4}x_{1}^{6}+34560b_{2}^{3}b_{4}x_{1}^{7}+51840b_{2}^{2}b_{4}x_{1}^{8}+38880b_{2}b_{4}x_{1}^{9}+$
$\quad 11664b_{4}x_{1}^{10}-192b_{2}^{5}x_{1}^{4}y_{1}^{2}-1728b_{2}^{4}x_{1}^{5}y_{1}^{2}-6048b_{2}^{3}x_{1}^{6}y_{1}^{2}-10368b_{2}^{2}x_{1}^{7}y_{1}^{2}-$
$\quad 8748b_{2}x_{1}^{8}y_{1}^{2}-2916x_{1}^{9}y_{1}^{2}+15360b_{2}^{4}b_{4}^{2}x_{1}^{4}+92160b_{2}^{3}b_{4}^{2}x_{1}^{5}+207360b_{2}^{2}b_{4}^{2}x_{1}^{6}+$
$\quad 207360b_{2}b_{4}^{2}x_{1}^{7}+77760b_{4}^{2}x_{1}^{8}-3072b_{2}^{4}b_{4}x_{1}^{3}y_{1}^{2}-23040b_{2}^{3}b_{4}x_{1}^{4}y_{1}^{2}-62208b_{2}^{2}b_{4}x_{1}^{5}y_{1}^{2}-$
$\quad 72576b_{2}b_{4}x_{1}^{6}y_{1}^{2}-31104b_{4}x_{1}^{7}y_{1}^{2}+192b_{2}^{4}x_{1}^{2}y_{1}^{4}+1728b_{2}^{3}x_{1}^{3}y_{1}^{4}+5616b_{2}^{2}x_{1}^{4}y_{1}^{4}+$
$\quad 7776b_{2}x_{1}^{5}y_{1}^{4}+3888x_{1}^{6}y_{1}^{4}+81920b_{2}^{3}b_{4}^{3}x_{1}^{3}+368640b_{2}^{2}b_{4}^{3}x_{1}^{4}+552960b_{2}b_{4}^{3}x_{1}^{5}+$
$\quad 276480b_{4}^{3}x_{1}^{6}-18432b_{2}^{3}b_{4}^{2}x_{1}^{2}y_{1}^{2}-110592b_{2}^{2}b_{4}^{2}x_{1}^{3}y_{1}^{2}-207360b_{2}b_{4}^{2}x_{1}^{4}y_{1}^{2}-$
$\quad 124416b_{4}^{2}x_{1}^{5}y_{1}^{2}+1536b_{2}^{3}b_{4}x_{1}y_{1}^{4}+11520b_{2}^{2}b_{4}x_{1}^{2}y_{1}^{4}+27648b_{2}b_{4}x_{1}^{3}y_{1}^{4}+20736b_{4}x_{1}^{4}y_{1}^{4}-$
$\quad 64b_{2}^{3}y_{1}^{6}-576b_{2}^{2}x_{1}y_{1}^{6}-1744b_{2}x_{1}^{2}y_{1}^{6}-1744x_{1}^{3}y_{1}^{6}+245760b_{2}^{2}b_{4}^{4}x_{1}^{2}+737280b_{2}b_{4}^{4}x_{1}^{3}+$
$\quad 552960b_{4}^{4}x_{1}^{4}-49152b_{2}^{2}b_{4}^{3}x_{1}y_{1}^{2}-221184b_{2}b_{4}^{3}x_{1}^{2}y_{1}^{2}-221184b_{4}^{3}x_{1}^{3}y_{1}^{2}+$
$\quad 3072b_{2}^{2}b_{4}^{2}y_{1}^{4}+18432b_{2}b_{4}^{2}x_{1}y_{1}^{4}+27648b_{4}^{2}x_{1}^{2}y_{1}^{4}-128b_{4}x_{1}y_{1}^{6}+393216b_{2}b_{4}^{5}x_{1}+$
$\quad 589824b_{4}^{5}x_{1}^{2}-49152b_{2}b_{4}^{4}y_{1}^{2}-147456b_{4}^{4}x_{1}y_{1}^{2}-256b_{6}y_{1}^{6}+262144b_{4}^{6}\Big{)}$
It is a tedious, yet easy task to check that every summand involving either $b_{4}$ or $b_{6}$ have $v_{2}$–valuation at least $14$. So, the list of terms with $v_{2}$–valuation exactly $12$ and $13$ is
$v_{2}=12$ $v_{2}=13$
$64b_{2}^{6}x_{1}^{6}$ $576b_{2}^{5}x_{1}^{7}$
$2160b_{2}^{4}x_{1}^{8}$ $2916b_{2}x_{1}^{11}$
$4860b_{2}^{2}x_{1}^{10}$ $-1728b_{2}^{4}x_{1}^{5}y_{1}^{2}$
$729x_{1}^{12}$ $-6048b_{2}^{3}x_{1}^{6}y_{1}^{2}$
$-192b_{2}^{5}x_{1}^{4}y_{1}^{2}$ $-2916x_{1}^{9}y_{1}^{2}$
$-8748b_{2}x_{1}^{8}y_{1}^{2}$ $1728b_{2}^{3}x_{1}^{3}y_{1}^{4}$
$192b_{2}^{4}x_{1}^{2}y_{1}^{4}$ $-576b_{2}^{2}x_{1}y_{1}^{6}$
$5616b_{2}^{2}x_{1}^{4}y_{1}^{4}$ $-1744x_{1}^{3}y_{1}^{6}$
$-64b_{2}^{3}y_{1}^{6}$
$-1744b_{2}x_{1}^{2}y_{1}^{6}$
There are precisely $10$ terms with $v_{2}$–valuation $12$ and $8$ with $v_{2}$–valuation $13$. Note that if we add the second and third terms in the left column (that is, $v_{2}$–valuation $12$), we get
$v_{2}(2160b_{2}^{4}x_{1}^{8}+4860b_{2}^{2}x_{1}^{10})=10+v_{2}(9x_{1}^{2}+4b_{2}^{2})=12+v_{2}((3x_{1}/2)^{2}+b_{2}^{2}).$
As $23x_{1}/$ 2 and $b_{2}$ are both odd integers, $v_{2}((3x_{1}/2)^{2}+b_{2}^{2})=1$ and
$v_{2}(2160b_{2}^{4}x_{1}^{8}+4860b_{2}^{2}x_{1}^{10})=13.$
Then we can add $2160b_{2}^{4}x_{1}^{8}+4860b_{2}^{2}x_{1}^{10}$ to a $v_{2}$-valuation $13$ term, say $576b_{2}^{5}x_{1}^{7}$ and
$v_{2}(2160b_{2}^{4}x_{1}^{8}+4860b_{2}^{2}x_{1}^{10}+576b_{2}^{5}x_{1}^{7})\geq 14.$
Therefore the long equality above can be written as
[…]
Pairing six terms of $v_{2}$–valuation $13$ we can write
[…]
If the sum of the $8$ terms with $v_{2}$–valuation $12$ had in turn $v_{2}$–valuation strictly greater than $13$ we would have arrived to our contradiction. So, let us assume
1313 1313 $=$ $v_{2}\Big{(}64b_{2}^{6}x_{1}^{6}+729x_{1}^{12}-192b_{2}^{5}x_{1}^{4}y_{1}^{2}-8748b_{2}x_{1}^{8}y_{1}^{2}+192b_{2}^{4}x_{1}^{2}y_{1}^{4}+$
$\quad 5616b_{2}^{2}x_{1}^{4}y_{1}^{4}-64b_{2}^{3}y_{1}^{6}-1744b_{2}x_{1}^{2}y_{1}^{6}\Big{)}$
Now let us make the substitution
$y_{1}^{2}=x_{1}^{3}+b_{2}x_{1}^{2}+8b_{4}x_{1}+16b_{6}$
to get
1313 1313 $=$ $v_{2}\Big{(}3872b_{2}^{4}x_{1}^{8}+5936b_{2}^{3}x_{1}^{9}-8364b_{2}^{2}x_{1}^{1}0-10492b_{2}x_{1}^{1}1+729x_{1}^{1}2+46464b_{2}^{3}b_{4}x_{1}^{7}+$
$\quad 6144b_{2}^{2}b_{4}x_{1}^{8}-111840b_{2}b_{4}x_{1}^{9}-12288b_{2}^{3}b_{4}^{2}x_{1}^{5}+24576b_{2}^{2}b_{4}^{2}x_{1}^{6}+92928b_{2}^{3}b_{6}x_{1}^{6}-$
$\quad 334848b_{2}b_{4}^{2}x_{1}^{7}+12288b_{2}^{2}b_{6}x_{1}^{7}-223680b_{2}b_{6}x_{1}^{8}-32768b_{2}^{3}b_{4}^{3}x_{1}^{3}-49152b_{2}^{3}b_{4}b_{6}x_{1}^{4}-$
$\quad 892928b_{2}b_{4}^{3}x_{1}^{5}+98304b_{2}^{2}b_{4}b_{6}x_{1}^{5}-1339392b_{2}b_{4}b_{6}x_{1}^{6}-96608b_{2}^{3}b_{4}^{2}b_{6}x_{1}^{2}-$
$\quad 49152b_{2}^{3}b_{6}^{2}x_{1}^{3}-5357568b_{2}b_{4}^{2}b_{6}x_{1}^{4}+98304b_{2}^{2}b_{6}^{2}x_{1}^{4}-1339392b_{2}b_{6}^{2}x_{1}^{5}-$
$\quad 393216b_{2}^{3}b_{4}b_{6}^{2}x_{1}-10715136b_{2}b_{4}b_{6}^{2}x_{1}^{3}-262144b_{2}^{3}b_{6}^{3}-7143424b_{2}b_{6}^{3}x_{1}^{2}\Big{)}$
Now the list of elements with $v_{2}$–valuation $\leq 13$ is much smaller:
$v_{2}=12$ $v_{2}=13$
$-8364b_{2}^{2}x_{1}^{10}$ $3872b_{2}^{4}x_{1}^{8}$
$729x_{1}^{12}$ $5936b_{2}^{3}x_{1}^{9}$
$-10492b_{2}x_{1}^{11}$
It is enough to show that the sum of the terms on the left column has $v_{2}$–valuation exactly $13$, as then the valuation of the full sum would be, at least, $14$. So, notice that
$v_{2}(729x_{1}^{12}-8364b_{2}^{2}x_{1}^{10})=10+v_{2}(3^{6}x_{1}^{2}-2^{2}2091b_{2}^{2})=12+v_{2}((3^{3}x_{1}/2)^{2}-2091b_{2}^{2}).$
Recall that $23^{3}x_{1}/$ 2 is (again) and odd integer and so is $b_{2}$. But, in order to have that $v_{2}((3^{3}x_{1}/2)^{2}-2091b_{2}^{2})\geq 2$ we must have
$(3^{3}x_{1}/2)^{2}\equiv 2091b_{2}^{2}\equiv 3b_{2}^{2}\mod 4,$
which is impossible. This finishes the proof. ∎
Note that we cannot change the hypothesis of $2|d$ by $2$ ramifies over $K$, because the curve $415.a$ 4 has good reduction in $2$, its torsion over $\mathbb{Q}$ is $\mathcal{C}_{4}$ and over $\mathbb{Q}(\sqrt{3})$ (where $2$ ramifies) is $\mathcal{C}_{8}$.
@@
Let $E/\mathbb{Q}$ be an elliptic curve and $K/\mathbb{Q}$ a quadratic extension. If there exists $P\in E(K)[16]$ of order 16 such that $[2]P\in E(\mathbb{Q})[8]$, then $E$ has bad reduction in $2$.
@@
We go again by contradiction and we assume that $E$ has good reduction in $2$, i.e. $\Delta_{1}$ is an odd integer. In order to get a Tate normal form, we use again the previous coordinates changes. Since $Q$ is an order $8$ point, there exists $t\in\mathbb{Q}$ such that
@@
(see [4, Chapter 4, §$\S$ 4]). In this way, the discriminant of the Tate normal form
@@
Note that, since we have that $[2]P\in E_{\mathrm{tors}}(\mathbb{Q})$ is a point of order eight, then $E_{\mathrm{tors}}(\mathbb{Q})={\mathcal{C}}_{8}$ or ${\mathcal{C}}_{2}\times{\mathcal{C}}_{8}$. In the second case, there is no room for growth over a quadratic extension (cf. [3, Theorem 2]). So we have $E_{\mathrm{tors}}(\mathbb{Q})={\mathcal{C}}_{8}$.
Now, Lemma 16 of [3] tell us that there exists $r\in\mathbb{Q}$ such that $t=r^{2}/(r^{2}+1)$ and
@@
$-\dfrac{\overline{a}_{1}\overline{a}_{2}}{\overline{a}_{3}}=\dfrac{(q^{4}-p^{4})-2(pq)^{2}}{p^{2}(p^{2}+q^{2})}.$
@@
(16) $-\dfrac{\overline{a}_{2}^{3}}{\overline{a}_{3}^{2}}=b=\dfrac{q^{2}(q^{2}-p^{2})}{(p^{2}+q^{2})^{2}},$
@@
which implies that $v_{2}(y_{1})=2(1+v_{2}(p))$. In particular, it is an even number greater than or equal to $4$. Note that $(x_{1},y_{1})$ is a torsion point of the curve $E_{2}$, which has integer coefficients. With a change of variables that preserves the coordinate $y$ and the discriminant, we can transform the equation of $E_{2}$ into an equation of the form $y^{2}=x^{3}+Ax+B$, and apply Nagell-Lutz locally at $2$ (see for example [11], Theorem 3.4 of Chapter VII and the proof of Corollary 7.2 of Chapter VIII).
@@
There are no elliptic curves defined over $\mathbb{Q}$ with $\mathcal{C}_{8}\subset E_{\mathrm{tors}}(\mathbb{Q})$ and good reduction at $2$ such that $\mathcal{C}_{16}\subset E_{\mathrm{tors}}(K)$, for a quadratic extension $K$.
@@ 4. The prime $\ell=3$
In this diagram, the dotted inclusions correspond to those contained in $\mathrm{SL}_{2}(\mathbb{F}_{3})$. This information will be relevant, as $\overline{\rho}_{E,3}(G_{K})\subset\mathop{\mathrm{SL}}\nolimits_{2}(\mathbb{F}_{3})$ if and only if $K\supset\mathbb{Q}(\sqrt{-3})$, the cyclotomic extension of cubic roots of unity. Additionally, we have marked with the symbol ()($\bullet$)( ∙ ) the subgroups that fix a nontrivial element (as justified below).
@@
    for every $a,b$ such that $ab\neq 0$. In particular, $ax=x$, $by=y$, for $a=b=-1$, hence $x=y=0$.
@@
Now, let us examine the possibilities that lead to the growth of torsion. Note that, after choosing a suitable basis of $E[\ell]$, we can assume that $\overline{\rho}_{E,3}(G_{\mathbb{Q}})$ coincides with a subgroup in the list above. If we have a subgroup $H\subset\overline{\rho}_{E,3}(G_{\mathbb{Q}})$ without fixed points, then any subgroup of $\overline{\rho}_{E,3}(G_{\mathbb{Q}})$ that is conjugate (inside $\mathrm{GL}_{2}(\mathbb{F}_{3})$) to it will also satisfy that it does not have any fixed points. Therefore $\overline{\rho}_{E,3}(G_{K})$ cannot be conjugated (inside $\mathrm{GL}_{2}(\mathbb{F}_{3})$) to such an $H$. We now proceed to analyse each case.
@@
  • ${\mathcal{C}}_{3}\subset{\mathcal{C}}_{6}$, but as ${\mathcal{C}}_{6}\subseteq\mathop{\mathrm{SL}}\nolimits_{2}(\mathbb{F}_{3})$, it does not occur as a torsion group over $\mathbb{Q}$.
@@
If $E_{\mathrm{tors}}(\mathbb{Q})\simeq\mathcal{C}_{3}$, necessarily $\mathop{\mathrm{Gal}}\nolimits(\mathbb{Q}(E[3])/\mathbb{Q})\simeq H_{3,1},H_{1,1},\mathcal{C}_{3}$. Of these, the only subgroup that has $\{\mathrm{id}\}$ as a subgroup of index $2$ is $H_{1,1}$. Thus, we have the possibility $\mathop{\mathrm{Gal}}\nolimits(\mathbb{Q}(E[3])/\mathbb{Q})$ conjugate to $H_{1,1}$ and $\mathop{\mathrm{Gal}}\nolimits(\mathbb{Q}(E[3])/K)=\{\mathrm{id}\}$. As $\mathrm{id}\in\mathop{\mathrm{SL}}\nolimits_{2}(\mathbb{F}_{3})$, it must happen that $K=\mathbb{Q}(\sqrt{-3})$.
@@
    Curve $219.a$ 2. It has Galois group $\mathop{\mathrm{Gal}}\nolimits(\mathbb{Q}(E[3])/\mathbb{Q})\simeq H_{1,1}$, which has a fixed point. That is, the torsion over $\mathbb{Q}$ is $\mathcal{C}_{3}$. Over the field $\mathbb{Q}(\sqrt{-3})$, it has a trivial Galois group, i.e., the torsion is $\mathcal{C}_{3}\times\mathcal{C}_{3}$.
@@
    Curve $180.b$ 1. It has the Galois group $\mathop{\mathrm{Gal}}\nolimits(\mathbb{Q}(E[3])/\mathbb{Q})\simeq B(3)$, which has no fixed points. The torsion over $\mathbb{Q}$ is $\mathcal{C}_{2}$, but over the field $\mathbb{Q}(\sqrt{3})$, it has Galois group isomorphic to $H_{3,1}$; the torsion is $\mathcal{C}_{6}$.
@@
    Curve $150.b$ 1. It has Galois group $\mathop{\mathrm{Gal}}\nolimits(\mathbb{Q}(E[3])/\mathbb{Q})\simeq B(3)$, which has no fixed points. The torsion over $\mathbb{Q}$ is trivial. Over the field $\mathbb{Q}(\sqrt{-15})$, it has Galois group isomorphic to $H_{3,1}$; the torsion is $\mathcal{C}_{3}$. The curve $1176.a$ 1 is another example of this situation.
@@
    Curve $3175.b$ 3. This curve has Galois group $\mathop{\mathrm{Gal}}\nolimits(\mathbb{Q}(E[3])/\mathbb{Q})\simeq{\mathcal{C}}_{s}(3)$, which has no fixed points. The torsion over $\mathbb{Q}$ is trivial. Over the field $\mathbb{Q}(\sqrt{-15})$ it has Galois group isomorphic to $H_{1,1}$.
In addition, we can characterize the growth of the torsion when $E$ has good reduction at $3$ and the discriminant of $K$ is a multiple of $3$. First, we prove there are no curves whose torsion grows from $\mathcal{C}_{1}$ to $\mathcal{C}_{9}$ and have good reduction at $3$.
@@
Let $E/\mathbb{Q}$ be an elliptic curve such that $E_{\mathrm{tors}}(\mathbb{Q})$ is trivial and $E_{\mathrm{tors}}(K)=\mathcal{C}_{9}$, with $K$ a quadratic field. Then, $E$ has bad reduction at 3.

### v2

@@ A Tate normal form is an equation of the form
So, $v_{2}(x_{1})=1$ and $v_{2}(b_{2})=0$. The pair $(x_{1},y_{1})$ is a solution of the equation
$y^{2}=x^{3}+b_{2}x^{2}+8b_{4}x+16b_{6}.$
The corresponding point in the original equation, i.e.
$y^{2}+a_{1}xy+a_{3}y=x^{3}+a_{2}x^{2}+a_{4}x+a_{6},$
is
$\Big(\dfrac{x_{1}}{4},\dfrac{y_{1}-a_{1}x_{1}-4a_{3}}{8}\Big),$
which is a $4$-torsion point. So, a strong version of Nagell-Lutz Theorem [14, VIII.7, Theorem 7.1] implies that
$1=v_{2}\left(\dfrac{x_{1}}{4}\right)\geq 0,$
a contradiction. ∎
Note that we cannot change the hypothesis of $2|d$ by $2$ ramifies over $K$, because the curve $15.a4$ has good reduction in $2$, its torsion over $\mathbb{Q}$ is $\mathcal{C}_{4}$ and over $\mathbb{Q}(\sqrt{3})$ (where $2$ ramifies) is $\mathcal{C}_{8}$.
@@
Let $E/\mathbb{Q}$ be an elliptic curve and $K/\mathbb{Q}$ a quadratic extension. If there exists $P\in E(K)[16]$ of order 16, then $E$ has bad reduction in $2$.
@@
Because of [1, Theorem 1.1], any point $P\in E(K)$ of order $16$, satisfy that $Q=2P\in E(\mathbb{Q})$. We go again by contradiction and we assume that $E$ has good reduction in $2$, i.e. $\Delta_{1}$ is an odd integer. In order to get a Tate normal form, we use again the previous coordinates changes. Since $Q$ is an order $8$ point, there exists $t\in\mathbb{Q}$ such that
@@
(see [5, Chapter 4, $\S$ 4]). In this way, the discriminant of the Tate normal form
@@
Note that, since we have that $2P\in E_{\mathrm{tors}}(\mathbb{Q})$ is a point of order eight, then $E_{\mathrm{tors}}(\mathbb{Q})={\mathcal{C}}_{8}$ or ${\mathcal{C}}_{2}\times{\mathcal{C}}_{8}$. In the second case, there is no room for growth over a quadratic extension (cf. [4, Theorem 2]). So we have $E_{\mathrm{tors}}(\mathbb{Q})={\mathcal{C}}_{8}$.
Now, [4, Lemma 16] tell us that there exists $r\in\mathbb{Q}$ such that $t=r^{2}/(r^{2}+1)$ and
@@
$\dfrac{\overline{a}_{1}\overline{a}_{2}}{\overline{a}_{3}}=\dfrac{(q^{4}-p^{4})-2(pq)^{2}}{p^{2}(p^{2}+q^{2})}.$
@@
(16) $\dfrac{\overline{a}_{2}^{3}}{\overline{a}_{3}^{2}}=b=\dfrac{q^{2}(q^{2}-p^{2})}{(p^{2}+q^{2})^{2}},$
@@
which implies that $v_{2}(y_{1})=2(1+v_{2}(p))$. In particular, it is an even number greater than or equal to $4$. Note that $(x_{1},y_{1})$ is a torsion point of the curve $E_{2}$, which has integer coefficients. With a change of variables that preserves the coordinate $y$ and the discriminant, we can transform the equation of $E_{2}$ into an equation of the form $y^{2}=x^{3}+Ax+B$, and apply Nagell-Lutz locally at $2$ (see for example [14, VII, Theorem 3.4] and the proof of [14, VIII, Corollary 7.2]).
@@
There are no elliptic curves defined over a quadratic field $K$ with $\mathcal{C}_{16}\subset E_{\mathrm{tors}}(K)$ and good reduction at $2$.
Proof.
It is a consequence of the previous result and [1, Theorem 1.1]
Note that the results of this section, along with Proposition 3, prove Theorem 2.
@@ 4. The prime $\ell=3$
In this diagram, the dotted inclusions correspond to those contained in $\mathrm{SL}_{2}(\mathbb{F}_{3})$. This information will be relevant, as $\overline{\rho}_{E,3}(G_{K})\subset\mathop{\mathrm{SL}}\nolimits_{2}(\mathbb{F}_{3})$ if and only if $K\supset\mathbb{Q}(\sqrt{-3})$, the cyclotomic extension of cubic roots of unity (see e.g.[16, Section 1.3]). Additionally, we have marked with the symbol $(\bullet)$ the subgroups that fix a nontrivial element (as justified below).
@@
    for every $a,b$ such that $ab\neq 0$. In particular, $ax=x$, $by=y$, for a=b=1a=b=-1, hence $x=y=0$.
@@
Now, let us examine the possibilities that lead to the growth of torsion. Note that, after choosing a suitable basis of $E[\ell]$, we can assume that $\overline{\rho}_{E,3}(G_{\mathbb{Q}})$ coincides with a subgroup in the list above. If we have a subgroup $H\subset\overline{\rho}_{E,3}(G_{\mathbb{Q}})$ without fixed points, then any subgroup of $\overline{\rho}_{E,3}(G_{\mathbb{Q}})$ that is conjugate (inside $\mathrm{GL}_{2}(\mathbb{F}_{3})$) to it will also satisfy that it does not have any fixed points. Therefore $\overline{\rho}_{E,3}(G_{K})$ cannot be conjugate (inside $\mathrm{GL}_{2}(\mathbb{F}_{3})$) to such an $H$. We now proceed to analyse each case.
@@
  • ${\mathcal{C}}_{3}\subset{\mathcal{C}}_{6}$, but as ${\mathcal{C}}_{6}\subseteq\mathop{\mathrm{SL}}\nolimits_{2}(\mathbb{F}_{3})$, it does not occur as an image of the Galois representation.
@@
If $E_{\mathrm{tors}}(\mathbb{Q})[3]\simeq\mathcal{C}_{3}$, necessarily $\mathop{\mathrm{Gal}}\nolimits(\mathbb{Q}(E[3])/\mathbb{Q})\simeq H_{3,1},H_{1,1},\mathcal{C}_{3}$. Of these, the only subgroup that has $\{\mathrm{id}\}$ as a subgroup of index $2$ is $H_{1,1}$. Thus, we have the possibility $\mathop{\mathrm{Gal}}\nolimits(\mathbb{Q}(E[3])/\mathbb{Q})$ conjugate to $H_{1,1}$ and $\mathop{\mathrm{Gal}}\nolimits(\mathbb{Q}(E[3])/K)=\{\mathrm{id}\}$. As $\mathrm{id}\in\mathop{\mathrm{SL}}\nolimits_{2}(\mathbb{F}_{3})$, it must happen that $K=\mathbb{Q}(\sqrt{-3})$.
@@
    Curve $19.a2$. It has Galois group $\mathop{\mathrm{Gal}}\nolimits(\mathbb{Q}(E[3])/\mathbb{Q})\simeq H_{1,1}$, which has a fixed point. That is, the torsion over $\mathbb{Q}$ is $\mathcal{C}_{3}$. Over the field $\mathbb{Q}(\sqrt{-3})$, it has a trivial Galois group, i.e., the torsion is $\mathcal{C}_{3}\times\mathcal{C}_{3}$.
@@
    Curve $80.b1$. It has the Galois group $\mathop{\mathrm{Gal}}\nolimits(\mathbb{Q}(E[3])/\mathbb{Q})\simeq B(3)$, which has no fixed points. The torsion over $\mathbb{Q}$ is $\mathcal{C}_{2}$, but over the field $\mathbb{Q}(\sqrt{3})$, it has Galois group isomorphic to $H_{3,1}$; the torsion is $\mathcal{C}_{6}$.
@@
    Curve $50.b1$. It has Galois group $\mathop{\mathrm{Gal}}\nolimits(\mathbb{Q}(E[3])/\mathbb{Q})\simeq B(3)$, which has no fixed points. The torsion over $\mathbb{Q}$ is trivial. Over the field $\mathbb{Q}(\sqrt{-15})$, it has Galois group isomorphic to $H_{3,1}$; the torsion is $\mathcal{C}_{3}$. The curve $176.a1$ is another example of this situation.
@@
    Curve $175.b3$. This curve has Galois group $\mathop{\mathrm{Gal}}\nolimits(\mathbb{Q}(E[3])/\mathbb{Q})\simeq{\mathcal{C}}_{s}(3)$, which has no fixed points. The torsion over $\mathbb{Q}$ is trivial. Over the field $\mathbb{Q}(\sqrt{-15})$ it has Galois group isomorphic to $H_{1,1}$.
In addition, we can characterize the growth of the torsion when $E$ has good reduction at $3$ and the discriminant of $K$ is a multiple of $3$. First, we prove there are no curves whose torsion grows from $\mathcal{C}_{1}$ to $\mathcal{C}_{9}$ and which have good reduction at $3$.
@@
Let $E/\mathbb{Q}$ be an elliptic curve such that $E_{\mathrm{tors}}(\mathbb{Q})$ is trivial and $E_{\mathrm{tors}}(K)=\mathcal{C}_{9}$, with $K$ a quadratic field. Then $E$ has bad reduction at 3.